# A Survey on Intent-aware Recommender Systems

### v1

@@ 5.2. Future Directions
The work by Rafique et al. (Rafique et al., 2023) focuses on smart cities application scenarios for IARS. The main assumption of their work is that in the future increasingly more traces of user behavior will be available, stemming from sensors, mobiles, Internet-of-Things devices etc., and that these pieces of information can be used for building next-generation recommender systems. In the current literature, sensor information is so far only used in a few selected works, e.g., in (Lee and Lee, 2014; Changmai et al., 2019). In addition to external sensor information, fine-grained user interaction data recorded by applications may also become frequently used in the future. Shi et al. (Shi and Ghedira, 2017), for example, used the user’s mouse scroll speed as a predictor variable in their IARS. Besides more knowledge about user behavior, also more information about the users’ situation and context, which may influence their intents, will become available through additional sensors or external knowledge sources. In addition, a further research avenue lies in leveraging the ‘world knowledge’ encoded in Large Language Models for improved IARS, as proposed in (Sun et al., 2023).
@@
As indicated above, we see a strong potential in future works that aim at understanding the idiosyncrasies of individual application use cases, and explicitly considering potential user intents in a given setting. Ideally, such investigations should be supported either by an underlying theory, e.g., from psychology as in (He et al., 2014a; Wang et al., 2020b, 2022e), or by empirical studies or domain-specific analyses, as done, e.g., for the music and video domains in (Mehrotra et al., 2019; Bénédict et al., 2023a; Schäfer et al., 2013; Schedl et al., 2018; Volokhin and Agichtein, 2018b). A better understanding of these application-specific intents then serves as a basis for improved models for predicting the user’s intent from their observed interactions, see, e.g., (Cheng et al., 2017) using observational and survey data at Pinterest for improved intent prediction.
@@
Last but not least, we also envision important future directions in the context of the user experience of IARS. In practice, modern media streaming and e-commerce sites use multiple rows of recommendations to account for the various possible intents that users might have when arriving on the site. While multi-row user interfaces are common in practice, only a few works exist on this topic in academic research, e.g., (Pu and Chen, 2007; Jannach et al., 2021a; Starke et al., 2023). More research is thus required in this area, for instance which intents should be supported in a given application setting, or how rows’content should be personalized and ordered; see (Gomez-Uribe and Hunt, 2015) for a discussion of the topic at Netflix. Furthermore, with the advent of LLM-based chatbots like ChatGPT, end users will become more and more be accustomed to natural language based advice-giving systems, and will increasingly use them as interactive and conversational recommender systems. In such usage scenarios, it may be quite common that end users will explicitly express their intents and situational context, e.g., “I am planning to go for dinner with my friends after attending the Mets game on Sunday. Any recommendations?” A key challenge in future research will thus not only be to understand the expressed intent, to map it to suitable item recommendations, but even to provide a user interface that visualizes the assumed intent, see (Ruotsalo et al., 2014) for a related approach towards intent-modeling in interactive search.

### v3

@@ 5.2. Future Directions
The work by Rafique et al. (Rafique et al., 2023) focuses on smart cities application scenarios for IARS. The main assumption of their work is that in the future increasingly more traces of user behavior will be available, stemming from sensors, mobiles, Internet-of-Things devices etc., and that these pieces of information can be used for building next-generation recommender systems. In the current literature, sensor information is so far only used in a few selected works, e.g., in (Lee and Lee, 2014; Changmai et al., 2019). In addition to external sensor information, fine-grained user interaction data recorded by applications may also become frequently used in the future. Shi et al. (Shi and Ghedira, 2017) and Sulikowski et al. (Sulikowski et al., 2021), for example, used the user’s mouse scroll speed as a predictor variable in their IARS. Besides more knowledge about user behavior, also more information about the users’ situation and context, which may influence their intents, will become available through additional sensors or external knowledge sources. Finally, another direction lies in leveraging user-generated content such as text reviews provided by the user community, as proposed in (Lin et al., 2023).
Generally, research in IARS today is to some extent hampered by the limited availability of datasets that contain information or indications of possible user intents. Similar problems exist for other types of recommender systems, such as context-aware systems, price- and profit aware systems, or ones that consider multiple stakeholders (Adomavicius and Tuzhilin, 2015; De Biasio et al., 2023; Abdollahpouri et al., 2020). However, this should not shy us away from directing our research efforts towards these highly relevant types of recommender systems. Instead, we believe that a paradigmatic shift is needed in how we do research on recommender systems. Instead of developing increasingly more complex models on a small set of frequently-used datasets that mainly consist of user-item interactions, an increased level of academia-industry collaborations is required. As mentioned above, this would allow us both to evaluate new models ‘in the wild’ and to develop richer datasets for more realistic offline evaluations.
@@
As indicated above, we see a strong potential in future works that aim at understanding the idiosyncrasies of individual application use cases, and explicitly considering potential user intents in a given setting. Ideally, such investigations should be supported either by an underlying theory, e.g., from psychology as in (He et al., 2014b; Wang et al., 2020b, 2022e), or by empirical studies or domain-specific analyses, as done, e.g., for the music and video domains in (Mehrotra et al., 2019; Bénédict et al., 2023a; Schäfer et al., 2013; Schedl et al., 2018; Volokhin and Agichtein, 2018b). A better understanding of these application-specific intents then serves as a basis for improved models for predicting the user’s intent from their observed interactions, see, e.g., (Cheng et al., 2017) using observational and survey data at Pinterest for improved intent prediction.
@@
We also envision important future directions in the context of the user experience of IARS. In practice, modern media streaming and e-commerce sites use multiple rows of recommendations to account for the various possible intents that users might have when arriving on the site. While multi-row user interfaces are common in practice, only a few works exist on this topic in academic research, e.g., (Pu and Chen, 2007; Jannach et al., 2021a; Starke et al., 2023). More research is thus required in this area, for instance which intents should be supported in a given application setting, or how rows’content should be personalized and ordered; see (Gomez-Uribe and Hunt, 2015) for a discussion of the topic at Netflix. Furthermore, with the advent of LLM-based chatbots like ChatGPT, end users will become more and more be accustomed to natural language based advice-giving systems, and will increasingly use them as interactive and conversational recommender systems. In such usage scenarios, it may be quite common that end users will explicitly express their intents and situational context, e.g., “I am planning to go for dinner with my friends after attending the Mets game on Sunday. Any recommendations?” A key challenge in future research will thus not only be to understand the expressed intent, to map it to suitable item recommendations, but even to provide a user interface that visualizes the assumed intent, see (Ruotsalo et al., 2014) for a related approach towards intent-modeling in interactive search.
Leveraging Large Language Models
The emergence of LLMs is disrupting research in recommender systems, and a rich variety of proposals to leverage LLMs for recommendation were put forward since the release of the first models that were based on the Generative Pre-Trained Transformer (GPT) architecture. Common ways of incorporating LLMs into the recommendation process include the direct usage of an AI assistant like ChatGPT, various ways of model fine-tuning, generating prompts and prompt tuning, as well as using the ‘world knowledge’ that is encoded in an LLM as a feature encoder, see (Wu et al., 204; Lin et al., 2024; Zhao et al., 2024). The effectiveness of different approaches of incorporating LLMs to improve sequential recommendation models was demonstrated in (Harte et al., 2023). Further, an iterative, prompt-based intent-aware technique for session-based problems was recently proposed by Sun et al. (Sun et al., 2024), leading to improved offline accuracy results. A notable aspect when using LLMs for intent-aware recommendation with the help of prompts is that an LLM can provide explanations and its chain-of-thought regarding the assumed current user intent. Given their rich knowledge, LLMs may also serve as a powerful tool to list possible user intents for a given recommender systems use case, thereby laying the foundation for future Explicit Intent Modeling approaches.